[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 04, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed May 4 12:49:32 UTC 2016


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
1. Scatterometer data flow from NESDIS to NAWIPS returned to normal Tuesday.

2. TSB added two models to AWIPS2. The GFS ensemble means and the GFS  
analysis run.

3. Mello added a list a key words to the collaboration server to trigger  
alarms at the request of the forecasters.


HSU:
1. TSB/NCO swapped out a bad mouse on Sawfish Tuesday.
Other items:
1. The HSW switch for AWIPS2-NHCN has a bad port (px2--> port8). TSB/NCO  
moved the px2 Ethernet cable to port 15.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160504/0ead8ff8/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list